Integrated Nanosystems Research Facility

  • The annual maintenance on the DI water systems for both INRF and BiON was completed. The maintenance consisted of treating the system with chlorine and an inhibitor to remove any bacteria. We confirmed the resistivity was back to 17 megohm in INRF but we still need to verify the BiON. The BiON should be useable Friday.
  • The STS is experiencing chill water issues for the dry mechanical pump. The pump tripped out on high temperature. We need to disassemble the chill water line to decontaminate to get operating temperatures back to normal. It should be ready by Monday.
  • E-beam evaporator: The high vac was disassembled, cleaned and O rings replace. High vac valve is opening and closing normally. We just received the chamber shrouding from the cleaning company. We need to finish cleaning the shutter actuator, reassemble, then re-gen the cryo pump. Possibly ready by Tuesday.
  • XeF2 pulse etcher chamber has been cleaned. We need to reconnect the pumping line to the chamber and verify the system is operating within normal parameters. It should be ready by Friday.
  • ION milling: We are verifying interconnections in the gun. We’ve identified one bad connection leading to the chamber feed through connector. The connector needs to be repaired.
  • Oriel flood exposure still working on high voltage power supply.
  • MA56 mercury lamp needs to be replaced and should be ready by Friday.
  • AB&M flood exposure mercury lamp needs to be replaced. We expect it to be back online by Friday.

INRF Staff